# VideoLLaMA 3: Frontier Multimodal Foundation Models for Image and Video Understanding

### v3

@@ Abstract
In this paper, we propose VideoLLaMA 3, a more advanced multimodal foundation model for image and video understanding. The core design philosophy of VideoLLaMA3 is vision-centric. The meaning of “vision-centric” is two-fold: the vision-centric training paradigm and vision-centric framework design. The key insight of our vision-centric training paradigm is that high-quality image-text data is crucial for both image and video understanding. Instead of preparing massive video-text datasets, we focus on constructing large-scale, high-quality image-text datasets. VideoLLaMA3 has four training stages: 1) Vision Encoder Adaptation, which enables the vision encoder to accept images of variable resolutions as input; 2) Vision-Language Alignment, which jointly tunes the vision encoder, projector, and LLM with large-scale image-text data covering multiple types (including scene images, documents, and charts) as well as text-only data. 3) Multi-task Fine-tuning, which incorporates image-text SFT data for downstream tasks and video-text data to establish a foundation for video understanding. 4) Video-centric Fine-tuning, which further improves the model’s capability in video understanding. As for the framework design, to better capture fine-grained details in images, the pretrained vision encoder is adapted to encode images of varying sizes into vision tokens with corresponding numbers, rather than a fixed number of tokens. For video inputs, we reduce the number of vision tokens according to their similarity so that the representation of videos will be more precise and compact. Benefiting from vision-centric designs, VideoLLaMA3 achieves compelling performances in both image and video understanding benchmarks.
@@ 1 Introduction
Recent years have witnessed the rapid growth of Large Language Models (LLMs) openai2024gpt4o ; TheC3 ; team2024gemini ; llama3 ; yang2024qwen2 ; liu2024deepseek , which significantly enhance natural language processing and understanding. The growth of LLMs enables intelligence at the language level. However, to progress further, we need intelligence that extends beyond language, as the world itself is inherently multimodal. Specifically, the model should be capable of perceiving both static scenes and dynamic environments, which necessitates the ability to understand images and videos. Building upon the success of LLMs, Multimodal LLMs (MLLMs) wang2024qwen2 ; aria ; wu2024deepseek ; chen2024expanding have been proposed for multimodal understanding.
@@
Existing MLLMs  li2023videochat ; lin2023video ; ataallah2024minigpt4 ; zhang2024longva ; zohar2024apollo ; song2024moviechat+ ; fei2024video ; liu2024kangaroo ; liu2024oryx ; wang2024internvideo2 ; jung2024pegasus ; Maaz2024VideoGPT+ ; cheng2024videollama ; sun2024video ; zhang2024video ; Maaz2023VideoChatGPT ; li2023mvbench ; chen2024sharegpt4video ; li2024llavaonevision ; mplug3 ; chen2023internvl ; longllava ; internlmxcomposer25 have made significant progress in multimodal understanding. Image-centric MLLMs wang2024qwen2 ; li2024llavaonevision ; internvl2.5 ; chen2023internvl ; molmo2024 ; liu2024nvila ; smolvlm2023 , leveraging high-quality image-text datasets li2024llava ; li2024llavaonevision ; tong2024cambrian ; mathew2021docvqadatasetvqadocument ; laurençon2024building ; masry2022chartqa ; huang-etal-2024-chart that are easier to collect and curate, have demonstrated strong performance in image understanding, such as visual question answering, OCR, and document understanding. Beyond static content such as images, video-centric MLLMs zhang2024video ; chen2024sharegpt4video ; cheng2024videollama ; damonlpsg2023videollama must tackle the added complexity of modeling the temporal dimension of videos, requiring models to handle dynamic content and capture dependencies across frames. This temporal complexity, combined with the need for large-scale video-text datasets that are often of lower quality and harder to annotate, makes video MLLMs more challenging. These challenges underscore the advantages of using image understanding as a foundation for video understanding. By extending the robust visual capabilities of image MLLMs, video models can focus on and better address the unique challenges of temporal and dynamic content modeling.
Inheriting from VideoLLaMA DBLP:conf/emnlp/ZhangLB23 and VideoLLaMA2 damonlpsg2024videollama2 , VideoLLaMA3, a more advanced multimodal foundation model, is proposed for image and video understanding. We design VideoLLaMA3 in a vision-centric way. Specifically, we propose a vision-centric training paradigm and vision-centric framework designs. For the training paradigm, considering the intrinsic relationship between image and video modalities - where videos are essentially sequences of temporally correlated images, we prioritize the improvement of image understanding, which in turn enhances the performance of video understanding. Moreover, compared to video-text data, image-text data is easier to collect and ensures higher data quality. For vision-centric framework designs, we propose adapting the vision encoder to handle images of any resolution during the image understanding enhancement stage and tuning the encoder to efficiently embed video inputs.
@@
On the model side, we enhance the vision encoder with two vision-centric designs: 1) we adapt the vision encoder to take images with dynamic resolutions as inputs, and 2) we lift the vision encoder to receive videos and compress the video tokens into more compacted representations. In previous methods li2024llavaonevision ; liu2023visualit ; liu2023improved ; internvl2.5 ; chen2023internvl , vision tokens are either with fixed numbers or with numbers among several fixed choices, which is an inflexible and unnatural way to represent images. To alleviate this limitation, we adapt the pretrained vision encoder to receive images with variable shapes. This is achieved by replacing the fixed positional embeddings with the Rotary Position Embedding (RoPE). We finetune the vision encoder in the vision encoder adaptation stage so that it can accommodate dynamic inputs. In this way, enabling it to process high-resolution images and images with unusual aspect ratios with minimal information loss. As for video inputs, we consider the redundant information in videos and propose to reduce the number of vision tokens to represent a video. The advantages of vision token compression are two-fold. One is to make the visual embeddings of videos more compact and precise so that the model can focus more on the dynamic parts of videos. The other is to save computation demands during training and inference for video understanding.
@@ 2.1 Any-resolution Vision Tokenization
In MLLMs, visual inputs are extracted into vision tokens for multimodal understanding. The common practice liu2023visualit ; liu2023improved is to extract visual inputs with a pre-trained ViT-based vision encoder. The pre-trained vision encoder only receives images with fixed resolutions, which introduces information loss. To alleviate information loss, AnyRes techniques li2024llavaonevision ; internvl2.5 ; chen2023internvl are proposed to split images into patches with fixed resolutions. Although AnyRes techniques increase the number of vision tokens, it is still inflexible and neglects the position relationship within an image when extracting vision tokens. In VideoLLaMA3, we adopt the idea of Any-resolution Vision Tokenization (AVT) dehghani2023patch ; wang2024qwen2 to dynamically process images and videos of any resolution. Concretely, we adapt the pre-trained vision encoder (ViT-based architectures) to handle variable resolutions by employing a strategy to replace the absolute position embeddings in ViT with 2D-RoPE su2024roformer . With AVT, images and videos of different resolutions are better represented with more details included in vision tokens. To make the vision encoder compatible with AVT, we fine-tune the vision encoder and the projector in the stage of Vision Encoder Adaptation (i.e., stage #1 in Figure 2) using scene images, document data, and scene images with texts.
@@ 2.2 Differential Frame Pruner
For videos, inputs usually have much more tokens than image inputs after tokenization. To reduce the computation demand for videos, we apply a per-frame 2$\times$2 spatial downsampling by bilinear interpolation to limit the context length within a certain range. Besides, considering that videos consist of frames with overlapping content, representing videos by stacking vision tokens from each frame leads to lengthy and redundant tokens. To further reduce the number of tokens of videos, we propose the Differential Frame Pruner (DiffFP) to prune the video tokens. Inspired by RLT (choudhury2024don, ), we compare the 1-norm distance between temporally consecutive patches within the pixel space. We consider temporally consecutive patches with smaller distances to be redundant, and the later patches can be pruned. Specifically, as shown in Figure 4, we first calculate the 1-norm distance between consecutive frames in the pixel space and then remove patches whose distances fall below a pre-defined threshold. Following RLT (choudhury2024don, ), we set the default threshold to 0.1.
@@ 2.3 Construction of High-Quality Image Re-Caption Dataset
To train our VideoLLaMA3, we constructed a high-quality image re-caption dataset, VL3-Syn7M. All images in this dataset are sourced from COYO-700M (kakaobrain2022coyo-700m, ) and processed using our proposed cleaning pipeline as below:
@@
5) Image Re-caption. After filtering and clustering the images, we proceed with detailed re-captioning. Brief captions are generated using InternVL2-8B chen2023internvl ; Chen2024HowFA , while the detailed captions are produced with InternVL2-26B chen2023internvl ; Chen2024HowFA . These two types of captions (VL3-Syn7M-short and VL3-Syn7-detailed) are employed at different stages of training to address varying needs.
@@ 3 Training
As illustrated in Figure 3, VideoLLaMA3 consists of four key components: a vision encoder, a video compressor, a projector, and a large language model (LLM). The vision encoder extracts visual tokens and is initialized with the pre-trained SigLIP (Zhai2023SigmoidLF, ). To reduce the number of vision tokens representing videos, a video compressor is employed. The projector bridges the features between the vision encoder and the LLM. For the LLM, we utilize Qwen2.5 models yang2024qwen2 .
Inspired by previous explorations in MLLMs wang2024qwen2 ; li2024llavaonevision ; zhang2024video , we develop video understanding capabilities based on strong image understanding foundations. To enable the model with strong image and video understanding capabilities simultaneously, the training of VideoLLaMA3 has four stages: 1) Vision Encoder Adaptation, 2) Vision-Language Alignment, 3) Multi-task Fine-tuning, and 4) Video-centric Fine-tuning. While the first three stages primarily focus on improving image understanding, the final stage is dedicated to video understanding. The details of the training stages are as follows:
1) Vision Encoder Adaptation. In this stage, we fine-tune the vision encoder, which is initialized with the pre-trained SigLIP (Zhai2023SigmoidLF, ), on a large-scale image dataset. During this stage, the vision encoder is made trainable, while the language decoder remains frozen. This fine-tuning transforms the encoder into a dynamic-resolution processor, enhancing its ability to process images of varying resolutions. Meanwhile, the projector is trained to better align the features of the vision encoder with those of the LLM.
@@ 3.1 Data Format
Streaming Video Sequence. For streaming video data, video and text tokens are interleaved in the sequence. Timestamps (i.e., “Time: xxs”) are inserted before the frame tokens, similar to video sequences. To mimic the interactive scenarios of streaming videos, Answer tokens (i.e., “GPT: xxx”) may appear within the sequence to denote contextualized outputs or interactions. The interleaved format ensures a seamless integration of video and textual data streams.
@@ 3.2 Data Mixture
Following the principle outlined in LLaVA-OneVision (li2024llavaonevision, ), i.e., “quality over quantity”, we conduct rigorous cleaning procedures to guarantee data quality. In this section, we provide a detailed description of the data mixture for each stage, as well as the synthesis and cleaning methods applied to different data subsets.
@@ 3.2.1 Vision Encoder Adaptation
  Scene Image VL3-Syn7M-short, LLaVA-Pretrain-558k (liu2023improvedllava, ), Objects365-Recap (Objects365, ), SA-1B-Recap (kirillov2023segment, ) 11.84M
Scene Text Image BLIP3-OCR-Recap (Xue2024xGenMMA, ) 0.93M
Document pdfa-eng-wds (pdfa, ), idl-wds idlwds 2.80M
The Vision Encoder Adaptation stage is designed to enhance the model’s ability to comprehend a wide range of diverse scenes and improve its feature extraction capacity, with a particular focus on capturing fine-grained information such as objects, regions, and text. As shown in Table 1, the training data in this stage combines scene images and document recognition images, along with a small portion of scene text images. It should be noted that all data labeled as "Recap" consists of captions generated with InternVL2-8B (chen2023internvl, ).
For scene images, our data sources include VL3-Syn7M-short, LLaVA-Pretrain-558K (liu2023improvedllava, ), Object365 (Objects365, ), and SA-1B (kirillov2023segment, ). Notably, the Object365 and SA-1B datasets are included to enhance data diversity, as images in this dataset are mainly complex scenes.
The scene text images are sourced from BLIP3-OCR (Xue2024xGenMMA, ). Both the brief recaption and the text content within the images are used as captions, and the text content caption following a left-to-right, top-to-bottom pattern across the image.
The document images used in this stage are a subset of pdfa-eng-wds (pdfa, ) and idl-wds (idlwds, ). A total of 2.8 million images were chosen from these two datasets, with the text content of the documents serving as image captions, following the reading order.
@@ 3.2.2 Vision-Language Alignment
  Scene Image VL3-Syn7M-detailed, Objects365-Recap (Objects365, ), SA-1B-Recap (kirillov2023segment, ), COCO2017-Recap (lin2014microsoft, ), ShareGPT4o (Chen2024HowFA, ), TextCaps (sidorov2020textcaps, ), ShareGPT4V (chen2023sharegpt4v, ), DenseFusion (li2024densefusion, ), LLaVA-ReCap (LCS-558K) (li2024llavaonevision, ) 12.56M
Scene Text Image Laion-OCR (schuhmann2022laion, ), COCO-Text (veit2016coco, ), TextOCR (singh2021textocr, ), BLIP3-OCR-Recap (Xue2024xGenMMA, ), LSVT (Sun2019ICDAR2C, ), ReCTS (liu2019icdar, ) 4.69M
Document SynthDoG-EN (kim2022ocr, ), SynthDoG-ZH (kim2022ocr, ), UReader-TR (Ye2023UReaderUO, ), FUNSD (funsd, ), DUDE (van2023icdar, ), Vary-600k (wei2023vary, ), pdfa-eng-wds (pdfa, ), idl-wds (idlwds, ) 2.68M
Chart Chart-to-Text (2022Chart, ) 0.04M
Fine-grained Osprey-724K (yuan2024osprey, ), MDVP-Data (lin2024draw, ), ADE20K-Recap (zhou2019semantic, ), Object365 (Objects365, ), Flickr-30K (young2014image, ), GranD (rasheed2024glamm, ) 1.00M
Text-only Evol-Instruct-143K (chen2024allava, ), Infinity-Instruct-code (InfinityInstruct2024, ), Infinity-Instruct-commonsense InfinityInstruct2024 , Infinity-Instruct-math (InfinityInstruct2024, ) 6.25M
@@
For scene images, we include COCO-2017 (veit2016coco, ), Object365 (Objects365, ), SA-1B (kirillov2023segment, ), ShareGPT4o (Chen2024HowFA, ), ShareGPT4V (chen2023sharegpt4v, ), DenseFusion (li2024densefusion, ), and LLaVA-Recap (LCS-558K) (li2024llavaonevision, ). For Object365, COCO-2017, and SA-1B datasets, we combined the original image annotations with InternVL2-26B chen2023internvl to recaption and generate detailed image captions.
The scene text images include a diverse set of Chinese and English scene text recognition datasets. These datasets, such as BLIP3-OCR (Xue2024xGenMMA, ), COCO-Text (veit2016coco, ), TextOCR (singh2021textocr, ), LSVT (Sun2019ICDAR2C, ), and ReCTS (liu2019icdar, ), provide varied examples of text in real-world environments. Furthermore, we filter images from the LAION dataset (schuhmann2022laion, ) to include those with clear and readable text, resulting in a collection of 3 million high-quality images, which we term as Laion-OCR dataset. For the Laion-OCR dataset captions, we include both the text content and the corresponding bounding box annotations of the text locations. The caption format is as follows:
@@
As for document images, we include pdfa-eng-wds (pdfa, ), idl-wds (idlwds, ), UReader-TR (Ye2023UReaderUO, ), Vary-600k (wei2023vary, ), and SynthDoG (kim2022ocr, ). SynthDoG dataset is constructed by generating synthetically accurate document images, avoiding human annotation errors and ensuring precise model training. Furthermore, we add the handwritten document dataset FUNSD (funsd, ) and the complex document dataset DUDE (van2023icdar, ). FUNSD provides annotated handwritten samples for handwriting recognition, while DUDE includes documents with complex layouts, enhancing the model’s ability to handle a variety of document types.
For chart images, since charts share many similarities with documents in terms of content presentation, we only include a limited amount of chart data. These data come from the Chart-to-Text (2022Chart, ) dataset.
For fine-grained images, we construct two types of data: region caption data and grounded caption data. Region caption data describes the content of specific regions within an image. These data are derived and constructed from the Ospery-724K (yuan2024osprey, ), Object365 (Objects365, ), ADE20K (lin2024draw, ), and MDVP-Data (zhou2019semantic, ) datasets. These data help the model to understand the details of the image at the region level. Grounded caption data consist of textual descriptions of objects with corresponding bounding box annotations, primarily constructed from the Flickr-30K (young2014image, ) and GranD (rasheed2024glamm, ) datasets. Both types of data enhance the model’s understanding of images, supporting more accurate object localization and recognition in complex scenes.

### v4

# ]DAMO Academy, Alibaba Group

@@ Abstract
In this paper, we propose VideoLLaMA 3, a more advanced multimodal foundation model for image and video understanding. The core design philosophy of VideoLLaMA3 is vision-centric. The meaning of “vision-centric” is two-fold: the vision-centric training paradigm and vision-centric framework design. The key insight of our vision-centric training paradigm is that high-quality image-text data is crucial for both image and video understanding. Instead of preparing massive video-text datasets, we focus on constructing large-scale, high-quality image-text datasets. VideoLLaMA3 has four training stages: 1) Vision Encoder Adaptation, which enables the vision encoder to accept images of variable resolutions as input; 2) Vision-Language Alignment, which jointly tunes the vision encoder, projector, and LLM with large-scale image-text data covering multiple types (including scene images, documents, and charts) as well as text-only data. 3) Multi-task Fine-tuning, which incorporates image-text SFT data for downstream tasks and video-text data to establish a foundation for video understanding. 4) Video-centric Fine-tuning, which further improves the model’s capability in video understanding. As for the framework design, to better capture fine-grained details in images, the pretrained vision encoder is adapted to encode images of varying sizes into vision tokens with corresponding numbers, rather than a fixed number of tokens. For video inputs, we reduce the number of vision tokens according to their similarity so that the representation of videos will be more precise and compact. Benefiting from vision-centric designs, VideoLLaMA3 achieves compelling performances in both image and video understanding benchmarks. Project repository: https://github.com/DAMO-NLP-SG/VideoLLaMA3.
@@ 1 Introduction
Recent years have witnessed the rapid growth of Large Language Models (LLMs) [1, 2, 3, 4, 5, 6], which significantly enhance natural language processing and understanding. The growth of LLMs enables intelligence at the language level. However, to progress further, we need intelligence that extends beyond language, as the world itself is inherently multimodal. Specifically, the model should be capable of perceiving both static scenes and dynamic environments, which necessitates the ability to understand images and videos. Building upon the success of LLMs, Multimodal LLMs (MLLMs) [7, 8, 9, 10] have been proposed for multimodal understanding.
@@
Existing MLLMs [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have made significant progress in multimodal understanding. Image-centric MLLMs [7, 29, 34, 31, 35, 36, 37], leveraging high-quality image-text datasets [38, 29, 39, 40, 41, 42, 43] that are easier to collect and curate, have demonstrated strong performance in image understanding, such as visual question answering, OCR, and document understanding. Beyond static content such as images, video-centric MLLMs [25, 28, 23, 44] must tackle the added complexity of modeling the temporal dimension of videos, requiring models to handle dynamic content and capture dependencies across frames. This temporal complexity, combined with the need for large-scale video-text datasets that are often of lower quality and harder to annotate, makes video MLLMs more challenging. These challenges underscore the advantages of using image understanding as a foundation for video understanding. By extending the robust visual capabilities of image MLLMs, video models can focus on and better address the unique challenges of temporal and dynamic content modeling.
Inheriting from VideoLLaMA [45] and VideoLLaMA2 [46], VideoLLaMA3, a more advanced multimodal foundation model, is proposed for image and video understanding. We design VideoLLaMA3 in a vision-centric way. Specifically, we propose a vision-centric training paradigm and vision-centric framework designs. For the training paradigm, considering the intrinsic relationship between image and video modalities - where videos are essentially sequences of temporally correlated images, we prioritize the improvement of image understanding, which in turn enhances the performance of video understanding. Moreover, compared to video-text data, image-text data is easier to collect and ensures higher data quality. For vision-centric framework designs, we propose adapting the vision encoder to handle images of any resolution during the image understanding enhancement stage and tuning the encoder to efficiently embed video inputs.
@@
On the model side, we enhance the vision encoder with two vision-centric designs: 1) we adapt the vision encoder to take images with dynamic resolutions as inputs, and 2) we lift the vision encoder to receive videos and compress the video tokens into more compacted representations. In previous methods [29, 47, 48, 34, 31], vision tokens are either with fixed numbers or with numbers among several fixed choices, which is an inflexible and unnatural way to represent images. To alleviate this limitation, we adapt the pretrained vision encoder to receive images with variable shapes. This is achieved by replacing the fixed positional embeddings with the Rotary Position Embedding (RoPE). We finetune the vision encoder in the vision encoder adaptation stage so that it can accommodate dynamic inputs. In this way, enabling it to process high-resolution images and images with unusual aspect ratios with minimal information loss. As for video inputs, we consider the redundant information in videos and propose to reduce the number of vision tokens to represent a video. The advantages of vision token compression are two-fold. One is to make the visual embeddings of videos more compact and precise so that the model can focus more on the dynamic parts of videos. The other is to save computation demands during training and inference for video understanding.
@@ 2.1 Any-resolution Vision Tokenization
In MLLMs, visual inputs are extracted into vision tokens for multimodal understanding. The common practice [47, 48] is to extract visual inputs with a pre-trained ViT-based vision encoder. The pre-trained vision encoder only receives images with fixed resolutions, which introduces information loss. To alleviate information loss, AnyRes techniques [29, 34, 31] are proposed to split images into patches with fixed resolutions. Although AnyRes techniques increase the number of vision tokens, it is still inflexible and neglects the position relationship within an image when extracting vision tokens. In VideoLLaMA3, we adopt the idea of Any-resolution Vision Tokenization (AVT) [49, 7] to dynamically process images and videos of any resolution. Concretely, we adapt the pre-trained vision encoder (ViT-based architectures) to handle variable resolutions by employing a strategy to replace the absolute position embeddings in ViT with 2D-RoPE [50]. With AVT, images and videos of different resolutions are better represented with more details included in vision tokens. To make the vision encoder compatible with AVT, we fine-tune the vision encoder and the projector in the stage of Vision Encoder Adaptation (i.e., stage #1 in Figure 2) using scene images, document data, and scene images with texts.
@@ 2.2 Differential Frame Pruner
For videos, inputs usually have much more tokens than image inputs after tokenization. To reduce the computation demand for videos, we apply a per-frame 2$\times$2 spatial downsampling by bilinear interpolation to limit the context length within a certain range. Besides, considering that videos consist of frames with overlapping content, representing videos by stacking vision tokens from each frame leads to lengthy and redundant tokens. To further reduce the number of tokens of videos, we propose the Differential Frame Pruner (DiffFP) to prune the video tokens. Inspired by RLT [51], we compare the 1-norm distance between temporally consecutive patches within the pixel space. We consider temporally consecutive patches with smaller distances to be redundant, and the later patches can be pruned. Specifically, as shown in Figure 4, we first calculate the 1-norm distance between consecutive frames in the pixel space and then remove patches whose distances fall below a pre-defined threshold. Following RLT [51], we set the default threshold to 0.1.
@@ 2.3 Construction of High-Quality Image Re-Caption Dataset
To train our VideoLLaMA3, we constructed a high-quality image re-caption dataset, VL3-Syn7M. All images in this dataset are sourced from COYO-700M [52] and processed using our proposed cleaning pipeline as below:
@@
5) Image Re-caption. After filtering and clustering the images, we proceed with detailed re-captioning. Brief captions are generated using InternVL2-8B [31, 53], while the detailed captions are produced with InternVL2-26B [31, 53]. These two types of captions (VL3-Syn7M-short and VL3-Syn7-detailed) are employed at different stages of training to address varying needs.
@@ 3 Training
As illustrated in Figure 3, VideoLLaMA3 consists of four key components: a vision encoder, a video compressor, a projector, and a large language model (LLM). The vision encoder extracts visual tokens and is initialized with the pre-trained SigLIP [54]. To reduce the number of vision tokens representing videos, a video compressor is employed. The projector bridges the features between the vision encoder and the LLM. For the LLM, we utilize Qwen2.5 models [5].
Inspired by previous explorations in MLLMs [7, 29, 25], we develop video understanding capabilities based on strong image understanding foundations. To enable the model with strong image and video understanding capabilities simultaneously, the training of VideoLLaMA3 has four stages: 1) Vision Encoder Adaptation, 2) Vision-Language Alignment, 3) Multi-task Fine-tuning, and 4) Video-centric Fine-tuning. While the first three stages primarily focus on improving image understanding, the final stage is dedicated to video understanding. The details of the training stages are as follows:
1) Vision Encoder Adaptation. In this stage, we fine-tune the vision encoder, which is initialized with the pre-trained SigLIP [54], on a large-scale image dataset. During this stage, the vision encoder is made trainable, while the language decoder remains frozen. This fine-tuning transforms the encoder into a dynamic-resolution processor, enhancing its ability to process images of varying resolutions. Meanwhile, the projector is trained to better align the features of the vision encoder with those of the LLM.
@@ 3.1 Data Format
Streaming Video Sequence. For streaming video data, video and text tokens are interleaved in the sequence. Timestamps (i.e. “Time: xxs”) are inserted before the frame tokens, similar to video sequences. To mimic the interactive scenarios of streaming videos, Answer tokens (i.e. “GPT: xxx”) may appear within the sequence to denote contextualized outputs or interactions. The interleaved format ensures a seamless integration of video and textual data streams.
@@ 3.2 Data Mixture
Following the principle outlined in LLaVA-OneVision [29], i.e. “quality over quantity”, we conduct rigorous cleaning procedures to guarantee data quality. In this section, we provide a detailed description of the data mixture for each stage, as well as the synthesis and cleaning methods applied to different data subsets.
@@ 3.2.1 Vision Encoder Adaptation
  Scene Image VL3-Syn7M-short, LLaVA-Pretrain-558k [55], Objects365-Recap [56], SA-1B-Recap [57] 11.84M
Scene Text Image BLIP3-OCR-Recap [58] 0.93M
Document pdfa-eng-wds [59], idl-wds [60] 2.80M
The Vision Encoder Adaptation stage is designed to enhance the model’s ability to comprehend a wide range of diverse scenes and improve its feature extraction capacity, with a particular focus on capturing fine-grained information such as objects, regions, and text. As shown in Table 1, the training data in this stage combines scene images and document recognition images, along with a small portion of scene text images. It should be noted that all data labeled as "Recap" consists of captions generated with InternVL2-8B [31].
For scene images, our data sources include VL3-Syn7M-short, LLaVA-Pretrain-558K [55], Object365 [56], and SA-1B [57]. Notably, the Object365 and SA-1B datasets are included to enhance data diversity, as images in this dataset are mainly complex scenes.
The scene text images are sourced from BLIP3-OCR [58]. Both the brief recaption and the text content within the images are used as captions, and the text content caption following a left-to-right, top-to-bottom pattern across the image.
The document images used in this stage are a subset of pdfa-eng-wds [59] and idl-wds [60]. A total of 2.8 million images were chosen from these two datasets, with the text content of the documents serving as image captions, following the reading order.
@@ 3.2.2 Vision-Language Alignment
  Scene Image VL3-Syn7M-detailed, Objects365-Recap [56], SA-1B-Recap [57], COCO2017-Recap [61], ShareGPT4o [53], TextCaps [62], ShareGPT4V [63], DenseFusion [64], LLaVA-ReCap (LCS-558K) [29] 12.56M
Scene Text Image Laion-OCR [65], COCO-Text [66], TextOCR [67], BLIP3-OCR-Recap [58], LSVT [68], ReCTS [69] 4.69M
Document SynthDoG-EN [70], SynthDoG-ZH [70], UReader-TR [71], FUNSD [72], DUDE [73], Vary-600k [74], pdfa-eng-wds [59], idl-wds [60] 2.68M
Chart Chart-to-Text [75] 0.04M
Fine-grained Osprey-724K [76], MDVP-Data [77], ADE20K-Recap [78], Object365 [56], Flickr-30K [79], GranD [80] 1.00M
Text-only Evol-Instruct-143K [81], Infinity-Instruct-code [82], Infinity-Instruct-commonsense [82], Infinity-Instruct-math [82] 6.25M
@@
For scene images, we include COCO-2017 [66], Object365 [56], SA-1B [57], ShareGPT4o [53], ShareGPT4V [63], DenseFusion [64], and LLaVA-Recap (LCS-558K) [29]. For Object365, COCO-2017, and SA-1B datasets, we combined the original image annotations with InternVL2-26B [31] to recaption and generate detailed image captions.
The scene text images include a diverse set of Chinese and English scene text recognition datasets. These datasets, such as BLIP3-OCR [58], COCO-Text [66], TextOCR [67], LSVT [68], and ReCTS [69], provide varied examples of text in real-world environments. Furthermore, we filter images from the LAION dataset [65] to include those with clear and readable text, resulting in a collection of 3 million high-quality images, which we term as Laion-OCR dataset. For the Laion-OCR dataset captions, we include both the text content and the corresponding bounding box annotations of the text locations. The caption format is as follows:
@@
As for document images, we include pdfa-eng-wds [59], idl-wds [60], UReader-TR [71], Vary-600k [74], and SynthDoG [70]. SynthDoG dataset is constructed by generating synthetically accurate document images, avoiding human annotation errors and ensuring precise model training. Furthermore, we add the handwritten document dataset FUNSD [72] and the complex document dataset DUDE [73]. FUNSD provides annotated handwritten samples for handwriting recognition, while DUDE includes documents with complex layouts, enhancing the model’s ability to handle a variety of document types.
For chart images, since charts share many similarities with documents in terms of content presentation, we only include a limited amount of chart data. These data come from the Chart-to-Text [75] dataset.
For fine-grained images, we construct two types of data: region caption data and grounded caption data. Region caption data describes the content of specific regions within an image. These data are derived and constructed from the Ospery-724K [76], Object365 [56], ADE20K [77], and MDVP-Data [78] datasets. These data help the model to understand the details of the image at the region level. Grounded caption data consist of textual descriptions of objects with corresponding bounding box annotations, primarily constructed from the Flickr-30K [79] and GranD [80] datasets. Both types of data enhance the model’s understanding of images, supporting more accurate object localization and recognition in complex scenes.